OpenStudy (sonyalee77):

A line passes through the point (0, –4) and has a slope of 3/2 . What is the equation of the line?

OpenStudy (sonyalee77):

@Savannah_Noelle

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(0\quad ,&-4) \end{array} \\\quad \\ slope = m= \cfrac{3}{2} \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)

OpenStudy (savannah_noelle):

@Sonyalee77 I'm not good at math, sorry!

OpenStudy (sonyalee77):

okay